Search

Date Entry File Actions
2024-12-20

Trailer chamber 5-3 and 5-6 overextended, repaired at 2PM (DW).

2024-12-20

Turned off program to trailer for Christmas break but will use this time to collect air samples to look at MDL further. Will break data up into small pieces for calculations.

2024-12-18

G1-5, G5-3, G5-4 chamber lids overextended. Repaired and adjusted pressure downwards on compressor to approx. 50psi

2024-12-09

Chamber 5-3, 5-4, 5-6, and 1-5 overextended. Repaired at 3PM kk

2024-12-02

Repaired G5-4, G5-5, G5-6, G5-3 chamber lids at 1:30PM.

2024-11-26

5-4 and 1-1 chambers needed repair.

2024-11-22

Repaired chamber 5-3 and 5-1. 5-6 and 5-4 were also over extended. KK

2024-10-30

Changed actuator w/ Dave on chamber 1-6, also found a line break on 1-4. 1-3. Repaired. 5-4 and 5-5 overextended lids.

2024-10-25

Adjusted all chamber heights in gashog to reflect reinstalled ht.

2024-10-24

Removed G5 autochambers for microplot harvest in the afternoon. Turned off chambers on gashog.

2024-10-23

Dave and Kevin replaced the chambers on G1 and restarted sampling. G1-3 overextending. Raised all tubing off the ground. Chambers removed 10/22 pm

2024-10-21

Dave removed chambers from G1 for the corn harvest at about 10:30 am.

2024-09-27

Chamber 5-3 lid was not attached. Repaired at approx 3:30PM KK

2024-09-24

Checked all chambers in G1 and G5, except G5-5 (closed at the time). Trimmed corn leaves out of the path of the closures on G1 chambers. Most G5s were in good shape and required minimal trimming on this visit. All G5 lid connections appeared to be intact and working properly.

2024-09-23

Trailer Licor stopped yesterday at 8 am. Reported a plugged inlet. I changed out the filters. No water visible in the input or output lines. The system was sampling 5-4. I reset the valves to standby and restarted the licor.

There is water in the NOx inlet line, but we are not running the NOx box right now.

2024-08-16

5-3 lid (one side in down position) repaired by DW. Restarted program at 3:36.

2024-08-13

Inlet plugged error on the Licor system. Reboot, started program. Looks like it was out overnight.

2024-08-11

The 750S router was rebooting randomly, resuling in late data errors from the gashog app. I’ve replaced the 750S router with the mango router I had there previously. However teh mango router does not have enough spare memory to do a good job of capturing packets.

2024-08-06

i replaced router and the ethernet cable between the router and the LI7810, to see if the licor will now stay on the network. I also enabled the wifi access point on the licor to hopefully have a connection to the licor if and when it goes off the network next.

2024-08-02

Licor analyzer went out on the trailer @ 3:30AM. Need to check and restart. Cooler on the NOx box is out.

2024-07-15

G5-3 chamber lid fixed. May have been okay up to point of checking lids, but arm attachment popped off when checking. Fixed.

2024-07-06

Licor stopped with a set point error. The cavity pressure was too low. I rebooted the licor, no change. When switched the valves to standby, the cavity pressure improved, but did not reach the set point. I’ve left the system in standby for now.

I disconnected the inlet and outlet tubes, no change. There was no visible moisture in the lines. Checked filters, and the top filter had a small area bent double. I replaced that filter, then the instrument indicated a plugged in let. I then replaced the bottom bluish filter with one of the white filter disks and then it appeared to start up normally.

I wonder if the bluish disks in the filter box are just impermeable supports?

I restared gashog sampling.

2024-07-03

replaced filters on the Licor in the BCSE trailer. I did not see any water, but the filters were pretty dirty.

2024-07-01

Clipped in chamber plants on automated chambers in G5.

2024-06-25

Lost power at the trailer for a good portion of the day. System back up and running by 4:30PM. MFC not reading properly but flow is steady at 370ml/min.

2024-06-24

Checked physical air flow at trailer mass flow controller. It is reading -84x ml/min, by reality is +372 ml/min. Checked outflow of MFC with Agilent flow meter.

2024-06-21

Turned on G1 auto chambers and reapplied chambers to gashog.

2024-06-20

Removing G1 lids for herbicide today. 4 Hr REI.

2024-06-17

Approximately 4pm, trimmed switchgrass and any other plant material growing out of the G5 automated chambers. Also checked lids and trimmed any plants outside of chambers that were interfering with lid seals.

2024-06-14

Chambers out and in the same day on G1 for sidedressing. Joe S helped me. Rain on 6/16 will promote fluxes. NOx box issues. Reconnected in the afternoon and Sven worked to collect the proper data stream. Running as of 6/17/24.

2024-06-12

Disassembled N2O LGR that had water contamination. Throughly cleaned both sides of the two reflective mirrors. Reassembled, will wait a day to try.

2024-06-06

Air filters on the licor were plugged. The first filter looks pretty dirty, the second filter looks fine. According to https://www.licor.com/env/support/Trace-Gas-Analyzers/topics/avoiding-water-ingress.html the filters are 0.2 um filters. Dried existing filters at 60C for 2 hours and re-installed. It no longer reports a plugged inlet.

Licor suggest the use of https://www.entegris.com/content/dam/shared-product-assets/wafergard-shared/datasheet-wafergard-gt-plus-and-gtl-in-line-5013.pdf to prevent water getting into the system.

2024-06-06

There was water in the normally open ambient input line. I hooked line up to the licor and Icos exhaust to clear it. Maybe the water did not enter through the sample lines. I can hook up a water tap to the ambient line since we don’t care about extra space in that line.

2024-06-05

the trailer sucked in water during the storm. The Licor reported a plugged inlet. I shut down the licor, disconnected both inlet and outlet lines from the licor and the icos. Both inlet and outlet lines had water droplets. I blew out the lines and left them disconnected to hopefully dry both instruments out. The ICOS reported very high H2O and very low N2O, but after about 30 minutes the H2O value dropped and the N2O value returned close to 300 ppb.

I will leave the lines diconnected for tonight and shut down the gashog app. We need to set up a water trap or some way of detecting water in the input lines and be able to divert or dry out the lines.

The inlet lines in the chambers were covered with droplets on the outside. I’m wondering if droplets that ran down the inlet tube in the chamber get sucked in if that chamber is sampled.

2024-05-30

It appears the ICOS needs to have the mirrors cleaned. Likely some water entered the system over the weekend. Caliubration worked but sensitivity is down, and noise up.

2024-05-29

Trailer went down over the weekend. ICOS concentrations way off and flow blocked. Found issue with valves, then recalibrated ICOS with our reference gas. Added lids back on G1 last Friday. Currently not running trailer until morning.

2024-05-20

Moved chamber 5-3 approximately 1 meter to the west to encompass some switchgrass plants, approximately 2pm. Trimmed all chambers in G5. Replaced clamp head on G1-5 (previous one had popped off a couple times).

2024-05-20

G1-6 showing negative CO2. Will check. Also removing chambers on SWFG for fertilization today.

2024-05-20

G5-3 updated chamber depths after moving: corner heights - 32.5cm, 32.2cm, 31.0cm, 30.0cm; center height - 29.5cm.

2024-05-15

BCSE automated chamber notes: G1-5: discovered a missing nut/bolt combo on lid that may leak air. Plugged temporarily with weatherstripping on both sides of hole. G5-1, 5-3, 5-4 chambers have very little to no switchgrass growth inside the chambers. May be time to move them to be more representative of the plot. G5-2: switchgrass growth to a few inches above top of chamber. This was trimmed back to being even/slightly below top of chamber.

2024-05-13

Re-installed G1 chambers after planting and liming management earlier in the day. Chambers all had to be shifted somewhat east to line up between newly planted rows. Four corner average heights recorded, along with one center measurement. (All numbers in centimeters.) G1-1: 30.7, 30.5, 29.7, 30.1cm, 30.5 center G1-3: 15.5, 15.6, 15.5, 15.6cm, 15.7cm center G1-2: 29.8, 30.1, 28.4, 28cm, 29cm center G1-6: 29.2, 28.9, 29.3, 29.5cm, 30cm center G1-5: 30, 29.8, 29.2, 28.9cm, 29.5cm center G1-4: 29.3, 29.5, 27.8, 28.2cm, 29.5cm center

2024-05-13

G1 chambers flagged and removed for planting and liming today. Chambers all removed by 9:30am. Will replace when farm management notifies us they have completed their work.

2024-04-29

G5-5 chamber had the arm disconnected, and one half lid was stuck closed. Reattached the arm.

2024-04-15

The Licor in the trailer disappeared from the network at 2024-4-12 19:30 EDT. I had to power down the instrument using the button on the side and restore power to get it to reconnect the network. Before restarting the Licor I tried to reconnect the network by removing and reinserting the ethernet cable, thinking that it would notice the loss of the ethernet cable and reset the network stack, but that did not work. After restarting the instrument indicated a plugged inlet, I removed both inlet and outlet lines and restarted the licor from the web panel. It started normally. I re-attached the inlet and outlet lines and restared the gashog app.

reduced the amount of blank samples measured.

2024-03-05

Replaced defective solenoid on G1-2 sampling line. The Licor was reading plugged inlet for this chamber, which suggests an issue downstream. We not only changed the solenoid but also the lid actuator which took almost a year to receive. Now have extra in stock.

2024-03-04

No NOx data is arriving since 2024-2-29 at 14:30. Checked connections and restarted the app. NOx data now showing.

2024-03-01

Dave reported that G1-2 remained closed for while he was at the trailer.

Licor reported a plugged inlet. It had been stuck sampling G1-2 since 2024-02-29 15:42. I disconnected the licor lines and rebooted. Waited for the laser to stabilize and reconnected the lines. I turned off G1-2 in case there is a plug in those lines.

2024-02-27

Dave checked chambers and NOx pump, all operating. Also Dave installed two TOMST soil moisture/temperature sensors on T2R1 near the installed CSI sensors reading soil moisture which were installed throughout the winter.

2024-02-26

Fixed disconnected lid on G5-3. Lid was in the closed position.

2024-02-14

Tracked down issue with G1-2 (finally). Both clamp actuators are leaking internally. Now have the Bimba actuators on hand (backordered for 10 months or more). Found them at Neff Engineering in Grand Rapids.

2024-02-05

Checked automated chambers. No issues seen with 1-6, though reports of issues from remote data. 5-4 was missing a clamp head. Found it and replaced it. May need a new one as a long-term fix.

2024-01-17

Disconnected the incoming and outgoing sampling lines from the Licor and ICOS in the trailer until we can clear the chambers and get them operational.

2024-01-11

G5-6, G1-3, G1-5 all overextended. Turned down compressor outgoing pressure to just over 52psi. Will shut down system today prior to snowstorm and incoming polar vortex.

2024-01-04

Reset G5-5 lid and reconnected clevis.

2024-01-02

G5-2. One lid is frozen to the base, the other was overextended. Moved the over-extended lid back but did not unfreeze the frozen lid.

2023-12-21

Found chamber G5-6 (control) had a hyperextended lid on the east side. Put back in place. Going back out to see if I can adjust it with a screwdriver to prevent it from recurring.

2023-12-08

LGR ICOS went down in the afternoon. No laser activity, mirror health warning, and high temp. Will swap with our last working instrument. Then decide which unit to send in for repair if any.

Notified Michigan Aerospace, they were on week two of collection with their instrument inline just before the return manifold.

2023-12-01

Valve on G6-6 leaking. Changed, but did not fix the issue. Block off outgoing tubing and recheck. One of the lid actuators at the box needs to be replaced.

2023-11-29

G5-3 has a leaky mechanical valve, lid. G1-3 has a frozen line on the lid and G1-6 may have a leak at the mechanical valve.

2023-11-20

Pressure leak on G5-5 over the weekend. Repaired and added oil to compressor. Pre-MI-aerspace instrument addition chamber settings. Must add one minute for extra volume of new laser instrument.

2023-11-06

Compressor was down on the trailer since 11/1 (I was on vacation). Repaired. Also installed the new pump for the NOx box and started the system.

2023-10-10

Looks like one or both clamps on G1-1 are leaking at chamber. Will need to order new parts.

2023-07-26

7/26 there was a problem with G5-2 lid solenoid on the trailer system. Found the signal wire to the switch was not connected, but also no lids were firing in close mode. Found a ground wire off as well. Reconnected the G5-2 lid wire and should all be working properly. DW checked the compressor as well.

2023-07-05

Trimmed switchgrass emerging above G5 automated chambers. Found G1-2 lids up, but north clamp down. May have blocked closure of lids over the extended weekend. Fixed north clamp.

KK- solenoid for this chamber not currently connected due to issue is either solenoid or clamps. Will address soon.

2023-06-30

G5-5 chamber east lid was not functioning from 6/18 - 6/29. Fluxes should be removed.

2023-06-27

BCSE G5, chamber 4 - west lid connector had become unattached. Removed old, rusty connector and replaced with new. (Approximately 2pm.)

2023-06-27

BCSE G5 automated flux chambers - trimmed all switchgrass in chambers to at/below chamber ceiling. Work done approximately 12:30pm.

2023-06-27

BCSE G1 automated flux chambers removed for fertilization approximately 10am. Returned to service by approximately 12:30pm.

2023-06-12

G1 chambers off for the last week. Turned on at 8:40AM today. No rain during this time frame. Was on vacation.

2023-05-25

Moved each G1 automated chamber ~8-12 inches east of previous location to place between corn rows. Planting offset from last year, requiring chamber movement to keep them between rows. All boxes moved between 10am and 11am.

2023-05-25

Dave reinstalled the G1 auto chambers in between corn rows. Corn emerging and chambers were set over new corn rows, now adjusted. New average heights added to gashog chambers for G1. Saw G5 fluxes increase after last Friday rain for a few days, now settling down. No rain since or in forecast.

2023-05-19

G5 sprayed with fertilizer yesterday. Not seeing much N2O response. I did inject N2O at input to ICOS, saw response. Also injected N2O into chamber gas port, also saw response.

2023-05-19

Removed lids from G1 chambers and removed two bases for potash application. Yesterday, returned G5 lids within an hour of fertilization (N). Very dry soil. Not seeing much N2O flux.

2023-05-19

3 G1 automated chambers that were moved eastward on 5/12 to accommodate the new sprayer were moved back to previous location (new sprayer didn’t work out, so old sprayer used). Chambers moved approximately 11am.

2023-05-12

G1 chambers out for corn planting (Dave). Dave reinstalled around 4PM. Moved a few chambers so herbicide spray could be completed without moving chambers over the weekend.

2023-05-11

At 9:20-9:30 AM used Ostrom Lab quantified N2O standard at 472ppb. Re-calibrated only the LGR ICOS instrument. ICOS CALIBRATION. System input and return lines reconnected about 11:15.

2023-05-02

Valve issue at 7:30AM. Restarting at 9:30AM. Trailer.

2023-04-11

Still using barriers on G1-4 and G5-6. Need to repair a few valves on the clamp and lid manifolds (leaking).

2023-04-06

NO fluxes are now noticeable for the first time in 2023. Warmer weather with high precipitation.

2023-04-05

No injections from 4-4-23 around 1PM to this morning. Sven restarted app.

2023-04-04

Returned G5 chambers around 2PM. Rained quite a bit over night.

2023-04-03

Removed G5 auto chambers around 10AM. Kept G1 chambers running.

2023-04-03

Found NOx box line cracked from pump exhaust to trap. Repaired. Was seeing very flat NOx readings for some time. The unit was not recirculating.

2023-03-28

Added a small amount of standard to the trailer ICOS to test response. All good. Will deploy N2O \”standard\” and Ultra zero standard this month. Then find a source for CH4 and CO2 standards in balance air.

2023-03-24

ICOS unit was hot and spectra flat. Highly negative spikes evident in trace. Turned the unit off for the weekend and sent spectra to ICOS (Bob). Monday 3/27 turned unit back on and shifted spectra back into place. Around 1030AM. Still working fine as of noon.

2023-02-01

Shut down compressor in trailer (per Kevin). Changed shut-off valves on manifolds to Swagelok. Checked chambers - all lids in order, all holding pressure except chamber 1-2.

2023-01-17

Licor at trailer crashed on Sunday night, then Monday afternoon (holiday). The cavity pressure is not pumping down to the given acceptable range of 40-45, reading 65kPa. Changing filters did not help. Licor claims there is a bad valve or a leak.

2023-01-12

Licor crashed around 9:45 or so. Restarted and then manually reloaded sampler and started run.

2023-01-12

eastern lid on chamber 5-5 was found closed on inspection at 8am. Arm attachment had come apart in the usual place. Reattached and arm/lid appear to be working correctly.

2023-01-11

Chamber lids hyperextending on 1-3,1-5, 5-3, and 5-6. Fixed 1-3, 5-3, and 5-6 - was able to reset and adjust arm/lid attachment to help minimize this in the future. 1-5 reset, but locking nuts could not be loosened to adjust arm. Kevin later adjusted the compressor pressure down to help prevent this from happening going forward.

2023-01-11

Dave mentioned actuator arm on 5-5 needed repair. He did so this morning. Yesterday I decreased the compressor pressure to 60 from 75 to prevent overextention of chamber lids. Seems to be working.

2023-01-06

After shutting down, restarted trailer yesterday mid-day. Checked chambers and cleared compressor before opening pressure to chambers. Clamp valve on G1-1 and 2 still needs troubleshooting. System running after Sven restarted program. Licor took over 30 minutes to restart after being shutdown for an extended period. Be patient! No snow on the ground currently, temps in the 30s.

2022-12-21

With holiday break and a major snowstorm rolling in. We decided to shutdown the automated trailer system until the new year. I turned off the compressor and the Licor instrument. Left the LGR ICOS and NOx cycling because they are harder to restart. Turned off around 12;40PM.

2022-12-20

Repaired trailer lid manifold valve 5 and 8. Several compressor changes caused particles to flow into manifold (yes there is a regulator and filter). Installed a Craftsman 30 Gal oil lubricated compressor purchased at Lowe\’s.

2022-12-19

Another compressor issue. Replaced oil-less compressor maybe two weeks ago. Thermal overload switches do not work on these compressors so not useful for us. Dave shutdown compressor at about 3PM. Dave also found several stainless tubes frozen. Will replace with Teflon over time.

2022-12-19

Many chamber SS inlets were plugged or partially plugged with ice - mostly cleaned out now. DW.

2022-12-16

Licor locked up several times over the last day. We checked each port separately and then added the sampler in. Nothing. Will continue to watch. We removed the NO measurement for now.

2022-12-15

Reset trailer in the afternoon. Valve 25 was \”stuck\” in the on position. When turned off, Licor plugged. Restarted. Also repaired G5-3 clevis.

2022-12-05

Trail compressor went down again. Need to move back to oil type compressors.

2022-12-05

changed compressor. automatic drain valve on compressor failed causing compressor to also fail.